Suppose a random sample of 36 is selected from a population with a standard deviation of 12. If the sample mean is 98, the 99% confidence Interval to estimate the population mean is what.
Answer #1

Sample size = n = 36

Sample mean = \bar{x} = 98

Population mean = \mu

Population standard deviation = \sigma = 12

100(1-\alpha)% confidence interval to estimate \mu is calculated as follows :-

\therefore 100(1-\alpha)% confidence interval to estimate \mu is given as :-  Vn

Given:-  100(1-\alpha)% = 99% \Rightarrow\alpha = 0.01 \Rightarrow\alpha/2 = 0.005

\therefore 99% confidence interval to estimate the population mean is given as :-  

  12 _ × 2.576| = 92.848, 103.152 12 2.576 , 98 + -Z0.005 | = 198__ × Vn t__Z 0.005, 7 + Vn

[The Z-scores are obtained from the Standard Normal table]